I need the awnser for an equation of a pool 10 feet long and 5 feet wide and it has 26 tiles around counting corners

Answer by rothauserc(4718)

The perimeter(P) of the pool is
P = 2*10 + 2*5 = 30 feet
each tile is 30/26 = 15/13 = 1.15 feet